A certain material has refractive indices 1.56, 1.60 and 1.68 for red, yellow and violet light respectively. (a) Calculate the dispersive power. (b) Find the angular dispersion produced by a thin prism of angle 6 made of this material.

Grade:upto college level

2 Answers

Deepak Patra
9 years ago
Given that Μ base r = 1.56, μ base y = 1.60, and μ base v = 1.68 (a) Dispersive power = ω = μ base v – μ base r/ μ base y – 1 = 1.68 – 1.56/1.60 – 1 =0.2 (b) Angular dispersion = (μ base v – μ base r)A = 0.12 × 6° = 7.2°
Kevin Nash
9 years ago
Sol. Given that Μ base r = 1.56, μ base y = 1.60, and μ base v = 1.68 (a) Dispersive power = ω = μ base v – μ base r/ μ base y – 1 = 1.68 – 1.56/1.60 – 1 =0.2 (b) Angular dispersion = (μ base v – μ base r)A = 0.12 × 6° = 7.2°